# Eliciting and Distinguishing Between Weak and Incomplete Preferences: Theory, Experiment and Computation

### v5

@@ Abstract
Recovering and distinguishing between the strict-preference, indifference and/or indecisiveness parts of a decision maker’s preferences is a challenging task but also important for testing theory and conducting welfare analysis. This paper contributes towards this goal by reporting on data from a lab experiment on riskless choice that were analyzed with novel theory-guided computational methods. The experiment included both Forced- and Free-Choice treatments. Its primary novelty consisted of allowing all subjects to select multiple alternatives at each menu. Based on a non-parametric goodness-of-fit criterion that we introduce, which generalizes intuitively a widely used pre-existing method to environments of multi-valued choices, each subjects’ decision data were tested against three structured general-choice models that feature maximization of stable but potentially weak and/or incomplete preferences. Nearly 60% of all subjects’ are well-explained by one of these models, typically with a unique model-optimal preference relation per subject. Importantly, preferences usually (80%) had a non-trivial indifference part and, where applicable, a clearly distinct indecisiveness part. The achieved uncoupling of revealed indifference and indecisiveness is documented empirically for the first time.
@@
Weak preferences; incomplete preferences; revealed indifference; revealed indecisiveness; choice correspondences; Jaccard-Houtman-Maks method; goodness-of-fit.
@@ 1         Introduction
The state in which a decision maker is indifferent between two or more choice alternatives plays a prominent role in many domains of economic analysis. As such, it is important to understand what kinds of observable decision environments and data-analytic methods can in principle allow for extracting an individual’s potentially weak preferences and distinguishing between their strict-preference and indifference parts. This is particularly pertinent if one also accepts that the individual in question may be indecisive, with preferences that are transitive but potentially incomplete. In this situation, it is not obvious how such observable decision data could be used to separate indifference from incomparability/indecisiveness.
These problems are important, both from a theoretical and a policy perspective. First, since many economic models of individual, collective or interactive decisions assume that agents have preferences with non-trivial indifference parts, elicitation of the agents’ weak preference relations would allow for testing these models’ descriptive relevance more accurately than if indifferences were assumed away. Additionally, knowing, for example, how many decision makers within a community consider e.g. a tree-planting program to be equally good to the development of a playground and how many of them have a strict preference instead would generally allow the community leader to arrive at a better decision than if all preferences were mistakenly interpreted to be strict. This example is made more concrete in the table below, which shows how eliciting single choices in the presence of indifferences may lead to a Pareto inefficient social outcome while eliciting multiple choices would not:
@@
True preferences $TreePlanting$ $TreePlanting$ $TreePlanting$
$\succ_{A}$ $\sim_{B}$ $\sim_{C}$ $TreePlanting$
$PlayGround$ $PlayGround$ $PlayGround$
Possible revealed preferences $TreePlanting$ $PlayGround$ $PlayGround$ $PlayGround$
@@
Revealed preferences when $TreePlanting$ $TreePlanting$ $TreePlanting$ $TreePlanting$
@@
This paper extends and combines recent developments in choice theory, computational revealed preference analysis and experimental economics in an ambitious attempt to elicit human decision makers’ potentially weak and/or incomplete preferences and, where relevant, to clearly separate their revealed strict preference, indifference and incomparability/indecisiveness components. The paper makes progress against this goal via a combination of methodological and empirical contributions. First, it proposes and reports on an implementation of a lab experimental design which elicits observable behavioural data that could be used to recover an individual’s weak preferences and distinguish between their strict-preference, indifference and, where relevant, incomparability/indecisiveness parts. The main feature of the design is that it allows subjects in both its Forced- and Free-Choice treatments to make multiple choices from every menu of alternatives they see by introducing incentives that help them make such possibly multi-valued choices in a preference-guided way. Subjects in the Free-Choice treatment can additionally avoid/delay making a choice from any menu at a small cost, without receiving any additional information about the alternatives. Subjects in the otherwise standard Forced-Choice treatment on the other hand are always required to choose at least one option. This experimental design was implemented in a lab environment of riskless choice from 50 distinct menus that were derived from 6 pairs of popular £10 gift cards, and data from 273 subjects are reported on. The primary novelty of this design compared to those in existing studies is that it enables the elicitation of multiple choices per subject in both its treatments. Evidently, this novelty is crucial for testing theories of choice that predict multi- and/or empty-valued choice correspondences as the outcome of preference maximization.
@@
The paper’s second methodological contribution consists of analysing these data by applying a novel computational method that is based on combinatorial optimization and allows for a model-rich approach towards recovering an individual’s potentially weak and/or incomplete preferences that pays particular attention to the fact that choices are possibly multi-valued. This method builds on and extends in the direction of multi-valued choice, and in a model-based way, the classic Houtman and Maks (1985) technique that is routinely used in empirical revealed preference tests of the rational choice/utility maximization hypothesis. More specifically, the original Houtman-Maks technique computes the maximal subset of a subject’s dataset that is consistent with rational choice, essentially allowing for an intuitive quantification of the subject’s behavioural proximity with that model. The extension that we use, which we refer to as the Jaccard-Houtman-Maks method, applies this approximation principle simultaneously to the model of (indifference-permitting) rational choice/utility maximization and two additional models of incomplete-preference maximization: (i) undominated choice, whereby an individual chooses the feasible alternative(s) that are not worse than anything else; (ii) dominant choice, whereby they choose the most preferred feasible alternative(s) if and only if those exist and defer otherwise.
Before justifying the addition of Jaccard’s name to those of the method’s original contributors we first note that, consistent with the questions that we raised earlier, all three models predict multi-valued choices under some of their admissible preference orderings, while the latter two also afford or predict two theoretical separations between revealed indifference and incomparability/indecisiveness. By finding which model is closest to explaining each subject’s behaviour, this method effectively allows for recovering –possibly approximately, and with the standard as if qualifications in place– both the individual’s deterministic decision rule and their preferences conditional on that rule. This explains the “model-based” part of the claimed extension. The addition of Jaccard’s name to the method we propose and use is due to the fact that, in order to account for the general multi-valuedness of subjects’ choices and of the model-predicted optimal choices, in this paper we compute the (dis-)similarity between these sets –which intuitively captures the proximity of observed and model-predicted choices– by means of the classic (and widely applied in several other fields) Jaccard (dis-)similarity metric between finite sets of discrete objects (Jaccard, 1901; Marczewski and Steinhaus, 1958; Levandovsky and Winter, 1971). This, in particular, allows one to have a more nuanced understanding of when a deviation of a subject’s set of chosen alternatives from the model-predicted ones is a “severe” mistake or a relatively “mild” one, contrary to the binary “pass/fail” approach of the straightforward extension of the HM method in this environment.
@@
Using the data collected from an experiment that was designed in the way described previously and analyzing them with the Jaccard-modified Houtman-Maks method outlined a few lines above, the paper makes several novel empirical contributions. First, despite the relatively large number of experimental decisions, 56% of all subjects are well-approximated (specifically, they are no more than 10% away from being perfectly explained) by one of the above three models of preference maximization. Furthermore, and perhaps surprisingly, the model-optimally recovered preferences of nearly 80% of all subjects in this group feature at least one indifference comparison between distinct alternatives. The prevalence of indifference that is suggested by this finding is important and merits further investigation in future studies. As far as the decomposition of preference-maximization types is concerned, 34% and 22% of all subjects are best matched by utility maximization and the two models of incomplete-preference maximization, respectively. In addition, a single optimal preference relation is recovered from 86% of all these subjects, highlighting the strengths of this method in sharp preference identification when applied on relatively rich –if incomplete– data sets and in relation to models that are in principle uniquely identifiable.
@@
The paper’s next empirical contribution –now from analysing the data at the aggregate level– is the documentation of a significant, and nearly identical, negative correlation between subjects’ choice consistency and their average response times. Although seemingly counter-intuitive, this interesting finding is broadly in line with a key prediction of the influential drift-diffusion neuro-economic model (see Ratcliff and McCoon, 2008, Alós-Ferrer et al., 2021 and references therein). The prediction might be interpreted as suggesting that shorter response times are more likely when the decision maker has a clear preference between the feasible options. Such clarity in turn would be expected to translate into more consistent active choices, which is indeed what we find. In addition, and despite the fact that subjects in our experiment were allowed to –and typically did– make multi-valued instead of single-valued choices, those in the Free-Choice treatment behaved significantly more consistently. This provides a positive robustness check of the main finding in the study by Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2022), which, unlike the present study, was predominantly designed to test for such potential differences in consistency, and did so with single-valued choice data. Finally, we exploit the richness and generality of the collected experimental data by also introducing and discussing the behaviour of new variables/statistics that are specific to multi-valued choice data, such as choice sizes and choice proportions, and analyse them relative to other relevant observables such as response times and choice-consistency indicators.
Although not constructed so as to enable a thorough investigation of such behaviours, we also note that utilizing two important aspects of the experimental design and its implementation allows for inferring that an additional 10% of the analysed subjects could be thought of as exhibiting a systematic satisficing behaviour (Simon, 1956; Caplin, Dean, and Martin, 2011; Reutskaja, Nagel, Camerer, and Rangel, 2011) or a preference for randomization (Agranov and Ortoleva, 2017, 2023; Dwenger, Kübler, and Weizsäcker, 2018; Cettolin and Riedl, 2019; Agranov, Healy, and Nielsen, 2022). Together with the model-based findings mentioned above, this increases the total proportion of behaviourally explainable subjects in our data to 66%. Finally, while the experiment was not designed so as to test for this either, we further investigated the potential presence of a “preference for flexibility” over menus (Kreps, 1979; Dekel, Lipman, and Rustichini, 2001). We did so by entertaining the possibility that some subjects might have “metavisualised” the actual experimental task of choosing potentially many alternatives from each menu, and responded to it by choosing from the collection of all menus that are derivable from that menu. Reassuringly, this analysis did not detect any notable patterns pointing towards a systematic such “metavisualised” preference for flexibility.
@@ 2.1         Three Deterministic Models of Preference-Maximizing Choice
The model of rational choice/utility maximization was characterized by Richter (1966) in a general environment that encompasses the one within which we are operating here. The model of undominated choice with incomplete preferences was, to the best of our knowledge, introduced by Schmeidler (1969) in a general equilibrium setting and was analyzed choice-theoretically under a variety of decision environments and preference structures by, most notably, Schwartz (1976), Bossert, Sprumont, and Suzumura (2005), Eliaz and Ok (2006), Bossert and Suzumura (2010) and Stoye (2015).111 This model is implicitly also used in stochastic-dominance applications of portfolio efficiency, along the lines studied in Levy (2016), Linton, Post, and Whang (2014) and Arvanitis, Scaillet, and Topaloglou (2023), for example. Dominant choice with incomplete preferences was studied theoretically in Gerasimou (2018a) [^2] A related pre-dating study is Dean (2008), which proposed a class of decision-avoidance models that focused on explaining the increasing prevalence of status quo bias -a phenomenon that is distinct from choice deferral- as menu size increases. When the decision problem contains no natural status quo and no dominant alternative, these models predict a compensatory decision process via which active choices -generally not fully consistent- are always made. Dominant choice with incomplete preferences on the other hand focuses on decision problems without a natural status quo option and features a non-compensatory decision process whereby the decision makermakes (fully consistent) active choices when and only when a most preferred alternative exists. and empirically in Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2022), with the latter analysis limited to an environment of single-valued non-forced choice experimental data. Those data, in particular, do not allow for testing the model of undominated choice or distinguishing between indifference and indecisiveness in either of the two models of choice with incomplete preferences, which is a primary goal of this paper.

### v6

@@ Abstract
Recovering and distinguishing between the strict-preference, indifference and/or indecisiveness parts of a decision maker’s preferences is a challenging task but also important for testing theory and conducting welfare analysis. This paper contributes towards this goal by reporting on data from a lab experiment on riskless choice that were analyzed with novel theory-guided computational methods. The experiment included both Forced- and Free-Choice treatments. Its main novelty consisted of allowing subjects to select multiple alternatives at each menu. Based on a new non-parametric goodness-of-fit criterion that we introduce, which generalizes a widely used pre-existing method to environments of multi-valued choices, each subject’s decisions were tested against three structured general choice models that feature maximization of stable but potentially weak and/or incomplete preferences. Nearly 60% of all subjects’ are well-explained by one of these models, typically with a unique model-optimal preference relation per subject. Importantly, revealed preferences typically have a non-trivial indifference part that, on average, accounts for up to 19% of all possible comparisons. In addition, 22% of all subjects are best explained by models of incomplete-preference maximization and reveal preferences that typically exhibit the distinctions between indifference and indecisiveness that these models afford or predict. These distinctions are documented empirically for the first time.
@@
Weak preferences; incomplete preferences; revealed indifference; revealed indecisiveness; choice correspondences; Jaccard-Houtman-Maks method; choice deferral.
@@ 1   Introduction
1.1   Motivation
The state in which a decision maker is indifferent between two or more choice alternatives in the sense of considering them to be equivalent plays a prominent role in many domains of economic analysis. As such, it is important to understand what kinds of observable decision environments and data-analytic methods can in principle allow for extracting an individual’s potentially weak preferences and distinguishing between their strict-preference and indifference parts. This is particularly pertinent if one also accepts that the individual in question may be indecisive, with preferences that are transitive but potentially incomplete. In this situation, it is not obvious how such observable decision data could be used to separate indifference from incomparability/indecisiveness.
These problems are important, both from a theoretical and a policy perspective. First, since many economic models of individual, collective or interactive decisions assume that agents have preferences with non-trivial indifference parts, elicitation of the agents’ weak preference relations would allow for testing these models’ descriptive relevance more accurately than if indifferences were assumed away. Additionally, knowing, for example, how many decision makers within a community consider e.g. a tree-planting program to be equally good to the development of a playground and how many of them have a strict preference instead would generally allow the community leader to arrive at a better decision than if all preferences were mistakenly interpreted to be strict. This example is made more concrete with the possibilities laid out in the table below, which shows how eliciting single choices in the presence of indifferences may lead to a Pareto inefficient social outcome while eliciting multiple choices would not:
@@ 1.1   Motivation
True preferences TreePlantingTreePlanting TreePlantingTreePlanting TreePlantingTreePlanting
$\succ_{A}$ $\sim_{B}$ $\sim_{C}$ TreePlantingTreePlanting
PlayGroundPlayGround PlayGroundPlayGround PlayGroundPlayGround
Possible revealed preferences TreePlantingTreePlanting PlayGroundPlayGround PlayGroundPlayGround PlayGroundPlayGround
@@
Revealed preferences when TreePlantingTreePlanting TreePlantingTreePlanting TreePlantingTreePlanting TreePlantingTreePlanting
@@
The relevance of understanding indecisiveness and, where necessary, distinguishing it from indifference/equivalence, also manifests itself in the the rapidly evolving landscape of Artificial-Intelligence (AI) Large Language Models (LLMs) that are currently in daily use by a sizeable and growing fraction of the global population. Specifically, in response to the persistent “hallucinations” problem [^1] This term refers to LLMs’ occasional answers to user questions that are confident- and authoritative-sounding but contain misleading, false or completely invented information. facing all LLMs that have become mainstream in the past few years, and which erodes public trust in these models, a segment of the machine-learning literature seeks to find ways in which this problem can be mitigated. Notably, in the very recent study that is co-authored by research scientists of a leading LLM provider, Kalai, Nachun, Vempala, and Zhang (2025) attribute the problem to the models’ incentives in their training and, especially, evaluation stages, which “reward guessing over acknowledging uncertainty”. Motivated by prior literature and pre-existing test-grading systems that penalize incorrect answers to dis-incentivize guessing but do not penalize “I don’t know”/abstaining answers, and where the relevant rules are known to test-takers in advance, the authors propose that LLM evaluation instructions at the models’ post-training stage explicitly state confidence targets within user prompts. For a confidence parameter $t$ between 0 and 1, for example, they suggest including the following at the end of every prompt: “Answer only if you are >t>t confident, since mistakes are penalized $\dfrac{t}{1-t}$ points, while correct answers receive 1 point, and an answer of “I don’t know” receives 0 points.” The potential enforcement of such a rule with a relatively high value of $t$ in an LLM’s post-training evaluation stage would clearly aim to prioritize accuracy and consistency over decisiveness, in partial analogy to Hitt’s (2019) analysis of decision-making in the US Supreme Court.
1.2   Contribution
@@
This paper extends and combines recent developments in choice theory, computational revealed preference analysis and experimental economics in an attempt to elicit human decision makers’ potentially weak and/or incomplete preferences and, where relevant, to separate their revealed strict preference, indifference and incomparability/indecisiveness components. The paper contributes in this direction with a combination of methodological and empirical innovations. First, it proposes and reports on an implementation of a lab experimental design which elicits observable behavioural data aiming at recovering an individual’s possibly weak preferences. The main feature of the design is that it allows subjects in both its Forced- and Free-Choice treatments to make multiple choices from every menu of alternatives they see by introducing incentives that help them make such possibly multi-valued choices in a preference-guided way. Subjects in the Free-Choice treatment can additionally avoid/delay making a choice from any menu at a small cost, without receiving any additional information about the alternatives. Subjects in the otherwise standard Forced-Choice treatment on the other hand are always required to choose at least one option. This experimental design was implemented in a lab environment of riskless choice from 50 distinct menus that were derived from 6 pairs of popular £10 gift cards, and data from 273 subjects are reported on. The main novelty of this design relative to existing ones is that it enables the elicitation of multiple choices per subject in both its treatments. Evidently, this is crucial for testing theories of choice that predict multi- and/or empty-valued choice correspondences as the outcome of preference maximization. While the design’s incentives are more closely aligned with testing the benchmark model of Rational Choice/Utility Maximization –importantly, allowing for possible indifferences– in this riskless environment, we also consider two distinct models of incomplete preference maximization to describe subjects whose behaviour is incompatible with that benchmark.
@@
Indeed, the paper’s second methodological contribution consists of analysing these data by applying a novel computational method that is based on combinatorial optimization and allows for a model-rich approach towards recovering an individual’s potentially weak and/or incomplete preferences that pays particular attention to the fact that choices are possibly multi-valued. This method builds on and extends in the direction of multi-valued choice –and in a model-based way– the classic Houtman and Maks (1985) technique that is routinely used in empirical revealed preference tests of the rational choice/utility maximization hypothesis. More specifically, the original Houtman-Maks technique computes the maximal subset of a subject’s dataset that is consistent with rational choice, essentially allowing for an intuitive quantification of the subject’s behavioural proximity with that model. The extension that we use, which we refer to as the Jaccard-Houtman-Maks method, accounts for the general multi-valuedness of choice data in these computations and applies this approximation principle simultaneously to the model of (indifference-permitting) rational choice/utility maximization and two additional models of incomplete-preference maximization: (i) undominated choice, whereby an individual chooses the feasible alternative(s) that are not worse than anything else; (ii) dominant choice, whereby they choose the most preferred feasible alternative(s) if and only if those exist and defer otherwise.
@@
Using the data collected from an experiment that was designed in the way described previously and analyzing them with the Jaccard-modified Houtman-Maks method outlined a few lines above, the paper makes several novel empirical contributions. First, despite the relatively large number of experimental decisions, 56% of all subjects are well-approximated (specifically, they are no more than 10% away from being perfectly explained) by one of the above three models of preference maximization. Furthermore, and perhaps surprisingly, the model-optimally recovered preferences of 77% of all subjects in this group feature at least one indifference comparison between distinct alternatives and, on average, almost 2.5 such comparisons, accounting for 15% of all possible binary comparisons between distinct alternatives. The prevalence of indifference suggested by these finding is important and merits further investigation in future studies. As far as the decomposition of preference-maximization types is concerned, 34% and 22% of all subjects are best matched by utility maximization and the two models of incomplete-preference maximization, respectively. In addition, a single optimal preference relation is recovered from 86% of all these subjects, highlighting the strengths of this method in sharp preference identification when applied on relatively rich –if incomplete– data sets and in relation to models that are in principle uniquely identifiable.
@@
The paper’s next empirical contribution, this time from analysing data at the aggregate level, is the documentation of a significant, and nearly identical, negative correlation between subjects’ choice consistency and their average response times. Although seemingly counter-intuitive, this interesting finding is broadly in line with a key prediction of the influential drift-diffusion neuro-economic model in binary forced-choice tasks (see Ratcliff and McCoon, 2008, Alós-Ferrer et al., 2021 and references therein). The prediction might be interpreted as suggesting that shorter response times are more likely when the decision maker has a clear preference between the feasible options. Such clarity in turn would be expected to translate into more consistent active choices, which is indeed what we find. In addition, and despite the fact that subjects in our experiment were allowed to –and typically did– make multi-valued instead of single-valued choices, those in the Free-Choice treatment behaved significantly more consistently. This provides a positive robustness check of the main finding in the study by Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2022), which, unlike the present study, was predominantly designed to test for such potential differences in consistency, and did so with single-valued choice data. Finally, the present study exploits the richness and generality of the collected experimental data by also introducing and discussing the behaviour of new variables/statistics that are specific to multi-valued choice data, such as choice sizes and choice proportions, and analyses them relative to other relevant observables such as response times and choice-consistency indicators.
1.3   Related Literature
The core of this paper’s experimental design extends in the multi-valued choice direction the design of Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2022). That study also reported on forced- and non-forced choice treatments but only elicited single-valued choices. This made it impossible to raise and answer the novel questions that this paper is mainly concerned with. Although a pre-publication version of that paper from 2016 had proposed a method of multi-valued choice construction, that method is considerably more restrictive than the one advanced here. In particular, contrary to this paper’s unrestricted elicitation of subjects’ multi-valued choices, that method used survey data at binary menus in order to augment with (constructed) transitive indifference relations the subjects’ single-valued choices at binary as well as non-binary menus. In work that followed the above-mentioned study, but which predates the present one, Bouacida (2021) proposed a different design that elicits multi-valued choices directly, by paying subjects a fixed extra amount for every additional alternative they choose from a menu of real-effort tasks, assigning them one of these tasks at random as a reward.
The design proposed herein differs from both these pre-existing ones and, in our view, materially so. Specifically, like Bouacida (2021) but unlike Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2016), we allow subjects to choose multiple alternatives at every menu directly, thereby avoiding the imposition of a priori consistency constraints on the collected multi-valued choice data. Yet, like Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2016) and unlike Bouacida (2021), we penalize inconsistent choices towards incentivizing subjects to reveal their true preferences. Furthermore, Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2016) implemented their design on all menus derived from a set of 5 headsets and with a 1-out-of-4 subject reward frequency [^2] A more complete list of the main differences between this experiment and those of that study is as follows: (i) (it) allows for potentially multi-valued choices vs requires necessarily single-valued ones; (ii) features 6 vs 5 choice alternatives (gift-card bundles vs headphones); (iii) presents 50 decisions vs 26; (iv) includes 273 subjects in the sample vs 161 & 121; (v) gives choice rewards to every subject vs 1 out of 4; (vi) makes no provision of any information about the options in the randomly selected menu before the 2nd decision vs provision of such information in the 1st experiment of the earlier study., while Bouacida (2021) implemented his on all menus derived from a set of 4 real-effort tasks where all subjects were rewarded. In addition, there are significant differences in the choice alternatives, number of choice problems, subjects’ reward frequency, and most importantly, in the data-analytic methods, research questions and conclusions of these papers.
Turning to the pre-existing literature on the elicitation of incomplete preferences, this has mainly revolved around choice over binary menus of lotteries and/or uncertain acts, and is reviewed in detail in Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2022). More summarily here: to identify incompleteness in such environments those studies have used choice-deferral alongside preference-for-flexibility models (Danan and Ziegelmeyer, 2006); partially incentivized methods of imprecise-preference revelation from menu lists (see Cubitt, Navarro-Martinez, and Starmer (2015) and references therein); and incentivized methods where incompleteness can potentially be revealed via certain patterns of randomized choice (Cettolin and Riedl, 2019; Agranov and Ortoleva, 2023). In general choice environments over durable goods on the other hand, strict incomplete preferences were documented in the non-forced-choice treatment of Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2022) via costly choice deferrals and restricted application of the model-based Houtman-Maks method to the dominant-choice model with strict incomplete preferences. By allowing for multi-valued choices in both Forced- and Free-Choice treatments, and by proposing and building the analysis upon the novel Jaccard-similarity modified variation of the Houtman-Maks method for multi-valued choice data, the design proposed in the present paper significantly improves upon Costa-Gomes et al. (2022) by enabling the analyst to test the hypothesis of of incomplete-preference maximization in the most general choice environments possible. Applied in these data, and based on the distinct choice-theoretic separations of these relations that were proposed in Eliaz and Ok (2006) and Gerasimou (2018a), the method allows to document empirically –and for the first time– the uncoupling of revealed indifference and incomparability/incompleteness that is predicted by these models.
@@ 2.1   Three Deterministic Models of Preference-Maximizing Choice
The model of rational choice/utility maximization was characterized by Richter (1966) in a general environment that encompasses the one within which we are operating here. The model of undominated choice with incomplete preferences was, to the best of our knowledge, introduced by Schmeidler (1969) in a general equilibrium setting and was analyzed choice-theoretically under a variety of decision environments and preference structures by, most notably, Schwartz (1976), Bossert, Sprumont, and Suzumura (2005), Eliaz and Ok (2006), Bossert and Suzumura (2010) and Stoye (2015).333 This model is implicitly also used in stochastic-dominance applications of portfolio efficiency, along the lines studied in Levy (2016), Linton, Post, and Whang (2014) and Arvanitis, Scaillet, and Topaloglou (2023), for example. Bouacida (2021) tests the no-indifference version of this model with the data collected under that study’s experimental design. Dominant choice with incomplete preferences was studied theoretically in Gerasimou (2018a) [^4] A related pre-dating study is Dean (2008), which proposed a class of decision-avoidance models that focused on explaining the increasing prevalence of status quo bias –a phenomenon that is distinct from choice deferral– as menu size increases. When the decision problem contains no natural status quo and no dominant alternative, these models predict a compensatory decision process via which active choices (not fully consistent in general) are always made. Dominant choice with incomplete preferences on the other hand focuses on decision problems without a natural status quo option and features a non-compensatory decision process whereby the agent makes (fully consistent) active choices when and only when a most preferred alternative exists. and empirically in Costa-Gomes, Cueva, Gerasimou, and Tejiščák (2022), with the latter analysis limited to an environment of single-valued non-forced choice experimental data. Those data, in particular, do not allow for testing the model of undominated choice or distinguishing between indifference and indecisiveness in either of the two models of choice with incomplete preferences, which is a primary goal of this paper.